P.19 differential equation 5x= 2y-x {y_x x) = 1 y)=0
Answer #1

Griven differential equation is X=24-X (2 H(o) = 1, y)=0 with initial conditions from ② Y lo = 400=1 and y - 2 - © using ②since, y (0)=0 and y (0) = 1. y (o)=0 => C, +(₂=0 Y(o)=1 =) -20,+. C = d. and C = + 1/3 ... y(t) = -be-2t + get and 2 (8)